a supermarket helper uses 1000 calories during an 8 hour work day.Suppose he takes a 15 minute break every three hours. Make some assumptions and calculate the power he delivers when he is working.
Answer #1

here,

the total energy used , E = 1000 Cal

time taken , t = 8 h - 2 * 15 min

t = 7.5 h

the power delivered when he is working, P = E /t

P = 1000 /7.5 Cal/h

P = 133.3 Cal/h = 133.3 * 4.184 /3600 W

P = 0.15 W

the power delivered when he is working is 0.15 W